OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (amistre64):

y-6 = -3(x--1) ; two -'s equals a positive ;) y-6 = -3(x+1)